Find the area between 𝑧 = 0 and each of the following.(used z-table).



1. z = 0.85



2. z = 1.27



3. z = 2.86



4. z = −1.05



5. z = −2.96

1
Expert's answer
2022-04-07T13:21:42-0400

"1.\ P(0<Z<0.85)=\\=P(Z<0.85)-P(Z<0)=\\=0.8023-0.5000=0.3023."

 ​

"2.\ P(0<Z<1.27)=\\=P(Z<1.27)-P(Z<0)=\\=0.8980-0.5000=0.3980."


"3.\ P(0<Z<2.86)=\\=P(Z<2.86)-P(Z<0)=\\=0.9979-0.5000=0.4979."


"4.\ P(-1.05<Z<0)=\\=P(Z<0)-P(Z<-1.05)=\\=0.5000-0.1469=0.3531."


"5.\ P(-2.96<Z<0)=\\=P(Z<0)-P(Z<-2.96)=\\=0.5000-0.0015=0.4985."